# Scale Where It Matters: Training-Free Localized Scaling for Diffusion Models

### v2

@@ G.3 Derived Corollaries
G.3.1 General Corollaries
@@
LoTTS achieves higher compute-normalized expected improvement than Global TTS if $\boxed{\begin{gathered}\frac{\rho}{C_{\ell}}\Big(q\overline{\delta}-(\tfrac{1}{\pi}-1)\,h_{\ell}\overline{\gamma}\Big)\\[6.0pt] >\;\frac{1}{C_{g}}\Big(\theta_{g}\overline{\delta}-(\tfrac{M}{s}-1)\,h_{g}\overline{\gamma}\Big)\end{gathered}}$
@@
Required recall.   Solving the inequality in Corollary 1 for $\rho$ yields
@@
Required precision.   A convenient sufficient condition for the denominator of equation G.3.1 to be positive is
$q\overline{\delta}-(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}\;>\;0\qquad\Longleftrightarrow\qquad\pi\;>\;\frac{1}{\,1+\frac{q\overline{\delta}}{h_{\ell}\overline{\gamma}}\,}.$ (22)
@@
G.3.2 Specialized Corollaries
We now present simplified dominance conditions obtained from Theorem 1 under practically relevant regimes. Each corollary corresponds directly to one of the scenarios in the Q&A of Section G.4.
Corollary 2 (Dominance Under Cost Asymmetry).
From Theorem 1, LoTTS dominates global TTS iff $\frac{\Delta_{\ell}^{\left(1\right)}}{C_{\ell}}\;>\;\frac{\Delta_{g}^{\left(1\right)}}{C_{g}}.$ (23) In the regime where $C_{\ell}\ll C_{g},\hskip 28.80008pt\Delta_{\ell}^{\left(1\right)}=\Theta(\Delta_{g}^{\left(1\right)})>0,$ condition equation 23 simplifies to $\boxed{\frac{\Delta_{\ell}^{\left(1\right)}}{C_{\ell}}\;\gg\;\frac{\Delta_{g}^{\left(1\right)}}{C_{g}}.}$ Thus LoTTS achieves strictly larger compute-normalized gain under any fixed compute budget $B$.
@@ G.3.2 Specialized Corollaries
$\mathbb{E}[\Delta r]_{\mathrm{local},B}=\frac{B}{C_{\ell}}\,\Delta_{\ell}^{\left(1\right)},\hskip 28.80008pt\mathbb{E}[\Delta r]_{\mathrm{global},B}=\frac{B}{C_{g}}\,\Delta_{g}^{\left(1\right)}.$
@@
Step 3: Cancel the common positive factor $B$.   Since $B>0$, inequality equation G.3.2 is equivalent to
@@
$C_{\ell}\ll C_{g},\hskip 28.80008pt\Delta_{\ell}^{\left(1\right)}\approx\Delta_{g}^{\left(1\right)}>0.$
The first assumption implies
@@
$\frac{\Delta_{\ell}^{\left(1\right)}}{C_{\ell}}=\Delta_{\ell}^{\left(1\right)}\frac{1}{C_{\ell}}\approx\Delta_{g}^{\left(1\right)}\frac{1}{C_{\ell}}\;\gg\;\Delta_{g}^{\left(1\right)}\frac{1}{C_{g}}=\frac{\Delta_{g}^{\left(1\right)}}{C_{g}}.$
Thus
@@
Step 6: Conclude dominance for any fixed budget $B$.   Substituting this back into equation G.3.2 shows that
@@
Assume $C_{\ell}=C_{g}$. Then Theorem 1 gives the exact condition $(\rho q-\theta_{g})\overline{\delta}\;>\;\Big(\tfrac{M}{s}-1\Big)h_{g}\overline{\gamma}+\rho\Big(\tfrac{1}{\pi}-1\Big)h_{\ell}\overline{\gamma}.$ (23) In the sparse-defect regime $s\ll M$ and with benign local edits $h_{\ell}\ll h_{g}$, the dominant terms yield the approximation $\boxed{\rho q\;\gtrsim\;\theta_{g}+\frac{M}{s}\,\frac{h_{g}\overline{\gamma}}{\overline{\delta}}.}$ Thus, under equal compute, dominance is governed by the relative scaling of the harm terms.
@@
$\frac{B}{C_{\ell}}\Big(\rho q\overline{\delta}-\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}\Big)\;>\;\frac{B}{C_{g}}\Big(\theta_{g}\overline{\delta}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\Big).$ (24)
Step 1: Specialize to $C_{\ell}=C_{g}$.   Under the equal-cost assumption $C_{\ell}=C_{g}$, the prefactor $B/C_{\ell}$ is strictly positive and common to both sides. Therefore we may cancel it from equation 24, yielding the exact dominance condition
$\rho q\overline{\delta}-\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}\;>\;\theta_{g}\overline{\delta}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}.$ (25)
Step 2: Collect all repair terms on the left and all harm terms on the right.   Rearranging equation 25 gives
$(\rho q-\theta_{g})\overline{\delta}$ $>\;(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\;+\;\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}.$ (26)
equation 26 is algebraically equivalent to the original dominance condition and valid without approximation.
Step 3: Apply sparse-defect assumption.   When $s\ll M$, we have
@@
so the global-harm term becomes
$(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\approx\frac{M}{s}h_{g}\overline{\gamma},$
which dominates all global contributions because it scales with the number of clean patches.
Step 4: Apply benign-local-edit assumption.   If $h_{\ell}\ll h_{g}$, then the local-harm term
$\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}$
is negligible compared to $\frac{M}{s}h_{g}\overline{\gamma}$, since (i) it does not scale with $M$, and (ii) $h_{\ell}\ll h_{g}$ suppresses its magnitude.
Thus, in the joint regime $(s\ll M)$ and $(h_{\ell}\ll h_{g})$, the right-hand side of equation 26 satisfies the asymptotic approximation
$(\tfrac{M}{s}-1)h_{g}\overline{\gamma}+\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}\;\approx\;\frac{M}{s}h_{g}\overline{\gamma}.$
Step 5: Divide both sides by $\overline{\delta}$.   Since $\overline{\delta}>0$, we divide equation 26 to obtain
$\rho q-\theta_{g}\;\gtrsim\;\frac{M}{s}\frac{h_{g}\overline{\gamma}}{\overline{\delta}},$
or equivalently
$\rho q\;\gtrsim\;\theta_{g}+\frac{M}{s}\frac{h_{g}\overline{\gamma}}{\overline{\delta}}.$
@@
For global Best-of-NN, $C_{g}(N)=N\,C_{g}\left(1\right),\hskip 28.80008pt\theta_{g}(N)=1-(1-\theta_{g}\left(1\right))^{N}.$ The compute-normalized improvement becomes $\boxed{\frac{1}{NC_{g}\left(1\right)}\Big(\theta_{g}(N)\overline{\delta}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\Big).}$ Since $\theta_{g}(N)$ increases sublinearly in NN while the harm term remains unchanged, global improvement saturates and cannot overcome the harm term by increasing $N$.
@@
Step 1: Define the Best-of-$N$ procedure and its cost.   A Best-of-$N$ trial draws $N$ independent global resampling samples and selects the best one. Since one global draw costs $C_{g}\left(1\right)$, performing $N$ draws costs
@@
Step 2: Compute the defect repair probability.   Let $\theta_{g}\left(1\right)$ be the probability that a single global draw repairs a defect. Under independence across draws, the probability that none of the $N$ draws repairs the defect is $(1-\theta_{g}\left(1\right))^{N}$. Thus the probability that at least one draw repairs it is
@@
Step 3: Substitute into the global gain expression.   Plugging $C_{g}(N)$ and $\theta_{g}(N)$ into the global expression in Theorem 1, the expected improvement under budget $B$ becomes
$\mathbb{E}[\Delta r]_{\mathrm{global},B}=\frac{B}{C_{g}(N)}\Big(\theta_{g}(N)\overline{\delta}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\Big).$
Step 4: Normalize by compute.   Dividing both sides by $B$ yields
$\frac{1}{C_{g}(N)}\Big(\theta_{g}(N)\overline{\delta}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\Big)$
$=$ $\frac{1}{NC_{g}\left(1\right)}\Big(\theta_{g}(N)\overline{\delta}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\Big),$
@@
Step 5: Show saturation as $N$ increases.   The function
$\theta_{g}\left(N\right)=1-\left(1-\theta_{g}\left(1\right)\right)^{N}$
is increasing and concave in $N$. Indeed, its incremental increase satisfies
@@
Step 6: Conclude saturation of compute-normalized global gain.   The compute-normalized gain contains the prefactor $1/(NC_{g}\left(1\right))$, which decreases linearly in $N$, while the harm term
@@
  • the factor $1/(NC_{g}\left(1\right))$ shrinks linearly.
@@
LoTTS fails to dominate whenever any one of the following conditions holds. These are exactly the scenarios where the dominance inequality in Theorem 1 reverses: Defects become dense ($s\!\approx\!M$), eliminating the sparse-defect advantage that normally suppresses global harm. Mask precision becomes extremely low, causing the false-positive harm term $\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}$ to dominate and outweigh the repair gains. Local repair probability is significantly smaller than the global one ($q\!\ll\!\theta_{g}$), reducing the expected improvement contributed by localized refinement.
@@
Step 2: Substitute the expressions from Theorem 1.   Using the explicit formulas for the expected improvements and cancelling the common positive factors yields
@@
Step 3: Group gain terms and harm terms.   Rearranging the inequality gives
$(\rho q-\theta_{g})\,\overline{\delta}\;\leq\;\Big(\tfrac{M}{s}-1\Big)h_{g}\overline{\gamma}+\rho\Big(\tfrac{1}{\pi}-1\Big)h_{\ell}\overline{\gamma}.$
Thus LoTTS fails precisely when the net local gain is no larger than the sum of the global harm and local false-positive harm.
@@
  • Extremely low precision ($\pi\ll 1$). If precision is very low, $(\tfrac{1}{\pi}-1)$ becomes large. Then the local harm term $\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}$ may dominate even when $h_{\ell}$ is small, causing LoTTS to apply many harmful edits to clean patches and making inequality equation G.3.2 hold.
@@
Step 5: Conclude the characterization of failure.   Each of the scenarios above corresponds to a way the right-hand side of equation G.3.2 can exceed the left-hand side: (i) global harm becomes small, (ii) local false-positive harm becomes large, (iii) local repair gain becomes too weak. These are precisely the regimes where
@@
Interpretation.   These corner cases correspond exactly to the settings in which LoTTS’s structural advantages—locality and sparse harm—no longer hold.

### v3

@@ G.3 Derived Corollaries
General Corollaries. The following corollary gives the general condition under which LoTTS achieves higher compute-normalized expected improvement than global TTS.
@@
LoTTS achieves higher compute-normalized expected improvement than Global TTS if and only if $\begin{gathered}\frac{\rho}{C_{\ell}}\!\left(q\,\overline{\delta}-(\tfrac{1}{\pi}-1)\,h_{\ell}\,\overline{\gamma}\right)\\[4.0pt] >\;\;\frac{1}{C_{g}}\!\left(\theta_{g}\,\overline{\delta}-(\tfrac{M}{s}-1)\,h_{g}\,\overline{\gamma}\right)\end{gathered}$
@@
Corollary 4 (Dominance Under Cost Asymmetry).
Let $\Delta_{\ell}^{(1)}$ and $\Delta_{g}^{(1)}$ denote the per-trial expected gains of localized and global resampling, respectively (i.e., the bracketed expressions in Theorem 2). From Theorem 2, LoTTS dominates global TTS iff: $\frac{\Delta_{\ell}^{\left(1\right)}}{C_{\ell}}\;>\;\frac{\Delta_{g}^{\left(1\right)}}{C_{g}}.$ (21) In the regime where $C_{\ell}\ll C_{g},\qquad\Delta_{\ell}^{\left(1\right)}=\Theta(\Delta_{g}^{\left(1\right)})>0,$ condition equation 21 simplifies to: $\frac{\Delta_{\ell}^{\left(1\right)}}{C_{\ell}}\;\gg\;\frac{\Delta_{g}^{\left(1\right)}}{C_{g}}.$ Thus LoTTS achieves strictly larger compute-normalized gain under any fixed compute budget $B$.
Required recall. Solving the inequality in Corollary 3 for $\rho$ yields:
@@
Required precision. A convenient sufficient condition for the denominator of Eq. equation G.3 to be positive is:
$q\overline{\delta}-(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}\;>\;0\;\;\Longleftrightarrow\;\;\pi\;>\;\frac{1}{1+\tfrac{q\overline{\delta}}{h_{\ell}\overline{\gamma}}}.$ (23)
@@
Specialized Corollaries. We now present simplified dominance conditions obtained from Theorem 2 under practically relevant regimes. Each corollary corresponds directly to one of the scenarios in the Q&A of Section G.4.
@@
$\mathbb{E}[\Delta r]_{\mathrm{local},B}=\frac{B}{C_{\ell}}\,\Delta_{\ell}^{\left(1\right)},\qquad\mathbb{E}[\Delta r]_{\mathrm{global},B}=\frac{B}{C_{g}}\,\Delta_{g}^{\left(1\right)}.$
@@
Step 3: Cancel the common positive factor $B$. Since $B>0$, inequality equation G.3 is equivalent to:
@@
$C_{\ell}\ll C_{g},\qquad\Delta_{\ell}^{\left(1\right)}\approx\Delta_{g}^{\left(1\right)}>0.$
The first assumption implies:
@@
$\frac{\Delta_{\ell}^{(1)}}{C_{\ell}}=\Delta_{\ell}^{(1)}\!\cdot\!\frac{1}{C_{\ell}}\approx\Delta_{g}^{(1)}\!\cdot\!\frac{1}{C_{\ell}}\;\gg\;\Delta_{g}^{(1)}\!\cdot\!\frac{1}{C_{g}}=\frac{\Delta_{g}^{(1)}}{C_{g}}.$
Thus:
@@
Step 6: Conclude dominance for any fixed budget $B$. Substituting this back into equation G.3 shows that:
@@
Assume $C_{\ell}=C_{g}$. Then Theorem 2 gives the exact condition $(\rho q-\theta_{g})\overline{\delta}\;>\;\rho\big(\tfrac{1}{\pi}-1\big)h_{\ell}\overline{\gamma}-\big(\tfrac{M}{s}-1\big)h_{g}\overline{\gamma}.$ (23) In the sparse-defect regime $s\ll M$ and with benign local edits $h_{\ell}\ll h_{g}$, the dominant terms yield the approximation $\rho q\;\gtrsim\;\theta_{g}-\frac{M}{s}\,\frac{h_{g}\overline{\gamma}}{\overline{\delta}}.$ Thus, under equal compute, dominance is governed by the relative scaling of the harm terms.
@@
$\frac{B}{C_{\ell}}\bigl(\rho q\overline{\delta}-\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}\bigr)$
$\quad>\;\frac{B}{C_{g}}\bigl(\theta_{g}\overline{\delta}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\bigr).$ (24)
Step 1: Specialize to $C_{\ell}=C_{g}$. Under the equal-cost assumption $C_{\ell}=C_{g}$, the prefactor $B/C_{\ell}$ is strictly positive and common to both sides. Therefore we may cancel it from equation G.3, yielding the exact dominance condition:
$\rho q\overline{\delta}-\rho(\tfrac{1}{\pi}\!-\!1)\,h_{\ell}\overline{\gamma}$
$\quad>\;\theta_{g}\overline{\delta}-(\tfrac{M}{s}\!-\!1)\,h_{g}\overline{\gamma}.$ (25)
Step 2: Collect all repair terms on the left and all harm terms on the right. Rearranging Eq. equation G.3 gives:
$(\rho q-\theta_{g})\overline{\delta}$ $>\;\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}\;-\;(\tfrac{M}{s}-1)h_{g}\overline{\gamma}.$ (26)
Eq. equation 26 is algebraically equivalent to the original dominance condition and valid without approximation.
Step 3: Apply sparse-defect assumption. When $s\ll M$, we have:
@@
so the global-harm reduction on the right-hand side becomes:
$(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\approx-\frac{M}{s}h_{g}\overline{\gamma},$
which is a large negative term that dominates the right-hand side because it scales with the number of clean patches.
Step 4: Apply benign-local-edit assumption. If $h_{\ell}\ll h_{g}$, then the local-harm term: $\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}$ is negligible compared to $\frac{M}{s}h_{g}\overline{\gamma}$, since (i) it does not scale with $M$, and (ii) $h_{\ell}\ll h_{g}$ suppresses its magnitude. Thus, in the joint regime $(s\ll M)$ and $(h_{\ell}\ll h_{g})$, the right-hand side of Eq. equation 26 satisfies the asymptotic approximation:
$\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\;\approx\;-\frac{M}{s}h_{g}\overline{\gamma}.$
Step 5: Divide both sides by $\overline{\delta}$. Since $\overline{\delta}>0$, we divide Eq. equation 26 to obtain:
$\rho q-\theta_{g}\;\gtrsim\;-\frac{M}{s}\frac{h_{g}\overline{\gamma}}{\overline{\delta}},$
or equivalently:
$\rho q\;\gtrsim\;\theta_{g}-\frac{M}{s}\frac{h_{g}\overline{\gamma}}{\overline{\delta}}.$
@@
We analyze Best-of-NN under an optimistic per-patch oracle model that independently selects the best outcome for each patch across NN draws. This upper-bounds the true Best-of-NN repair rate (which selects a single image), making the saturation conclusion conservative. For global Best-of-NN, $C_{g}(N)=N\,C_{g}(1),\qquad\theta_{g}(N)=1-\bigl(1-\theta_{g}(1)\bigr)^{\!N}.$ The compute-normalized improvement becomes $\frac{s}{NC_{g}\!\left(1\right)}\bigl(\theta_{g}(N)\overline{\delta}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\bigr).$ Since $\theta_{g}(N)$ increases sublinearly in NN while the harm term remains unchanged, global improvement saturates and cannot overcome the harm term by increasing $N$.
@@
Step 1: Define the Best-of-$N$ procedure and its cost. A Best-of-$N$ trial draws $N$ independent global resampling samples and selects the best one. Since one global draw costs $C_{g}\left(1\right)$, performing $N$ draws costs:
@@
Step 2: Compute the defect repair probability. Let $\theta_{g}\left(1\right)$ be the probability that a single global draw repairs a defect. Under independence across draws, the probability that none of the $N$ draws repairs the defect is $(1-\theta_{g}\left(1\right))^{N}$. Thus the probability that at least one draw repairs it is:
@@
Step 3: Substitute into the global gain expression. Plugging $C_{g}(N)$ and $\theta_{g}(N)$ into the global expression in Theorem 2, the expected improvement under budget $B$ becomes:
$\mathbb{E}[\Delta r]_{\mathrm{global},B}=\frac{Bs}{C_{g}(N)}\bigl(\theta_{g}(N)\overline{\delta}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\bigr).$
Step 4: Normalize by compute. Dividing both sides by $B$ yields:
$\frac{s}{C_{g}(N)}\bigl(\theta_{g}(N)\overline{\delta}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\bigr)$
$={}\;$ $\frac{s}{NC_{g}(1)}\bigl(\theta_{g}(N)\overline{\delta}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\bigr),$
@@
Step 5: Show saturation as $N$ increases. The function $\theta_{g}\left(N\right)=1-\left(1-\theta_{g}\left(1\right)\right)^{N}$ is increasing and concave in $N$. Indeed, its incremental increase satisfies:
@@
Step 6: Conclude saturation of compute-normalized global gain. The compute-normalized gain contains the prefactor $s/(NC_{g}\left(1\right))$, which decreases linearly in $N$, while the harm term
@@
  • the factor $s/(NC_{g}\left(1\right))$ shrinks linearly.
@@
Assume $C_{\ell}=C_{g}$. LoTTS fails to dominate whenever any one of the following conditions holds. These are exactly the scenarios where the dominance inequality in Theorem 2 reverses: Defects become dense ($s\!\approx\!M$), eliminating the sparse-defect advantage that normally suppresses global harm. Mask precision becomes extremely low, causing the false-positive harm term $\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}$ to dominate and outweigh the repair gains. Local repair probability is significantly smaller than the global one ($q\!\ll\!\theta_{g}$), reducing the expected improvement contributed by localized refinement.
@@
Step 2: Substitute the expressions from Theorem 2. Using the explicit formulas for the expected improvements and the assumption $C_{\ell}=C_{g}$, the common positive factor $Bs/C_{g}$ cancels, yielding:
@@
Step 3: Group gain terms and harm terms. Rearranging the inequality gives:
$(\rho q-\theta_{g})\,\overline{\delta}\;\leq\;\rho\big(\tfrac{1}{\pi}-1\big)h_{\ell}\overline{\gamma}-\big(\tfrac{M}{s}-1\big)h_{g}\overline{\gamma}.$
Thus LoTTS fails precisely when the net local repair advantage is no larger than the local false-positive harm minus the global harm.
@@
  • Extremely low precision ($\pi\ll 1$). If precision is very low, $(\tfrac{1}{\pi}-1)$ becomes large. Then the local harm term $\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}$ may dominate even when $h_{\ell}$ is small, causing LoTTS to apply many harmful edits to clean patches and making inequality equation G.3 hold.
@@
Step 5: Conclude the characterization of failure. Each of the scenarios above corresponds to a way the right-hand side of equation G.3 can exceed the left-hand side: (i) global harm becomes small, (ii) local false-positive harm becomes large, (iii) local repair gain becomes too weak. These are precisely the regimes where
@@
Interpretation. These corner cases correspond exactly to the settings in which LoTTS’s structural advantages, locality and sparse harm, no longer hold.